Answer just part b

) All firms in a perfectly competitive industry face the same long-run average cost curve, AC = 0.05q – 5 + 500/q, and the same long-run marginal cost curve given by MC = 0.1q – 5. The market demand for the product of these firms is QD = 100,000 – 10,000P. i. Calculate the equilibrium price and quantity. ii. Assuming the market is in long-run equilibrium, how many firms will be on the market? (b) Suppose the demand for cotton T-shirts is given by QD =1900–200P, where Q is the number of T-shirts and P is the price in dollars per T-shirt. The long-run supply curve for T-shirts is given by QS = 50P–100. Calculate consumer and producers surplus at equilibrium price and quantity.
Answer #1

QD = 1900 - 200P

QS = 50P - 100

At equilibrium QD = QS

1900 - 200P = 50P - 100

250P = 2000

P = 8

Q = 50*8 - 100 = 400-100 = 300

Finding Y intercept

When QD =0, P = 1900/200 = 9.5

When QS = 0, P = 100/50 = 2

Consumer surplus = 0.5 * (9.5-8)*300 = 150*1.5 = 225

Producer surplus = 0.5*(8-2)*300 = 150*6 = 900